a car is moving at a speed of 48 km/h accelerates 100 m up a steep hill, so that at the top of the hill its speed is 59 km/h. if the car's mass is 1100 kg, what is the magnitude of the net force acting on it?
1
Expert's answer
2018-11-16T10:40:10-0500

The acceleration

a=(v_f^2-v_i^2)/2s=(〖(59〗^2-〖48〗^2))/(2×100×〖3.6〗^2 )=0.45 m/s^2

The net force

F=ma=1100×0.45=499 N